OpenStudy (anonymous):

factor by grouping 3c^2+1c-24

OpenStudy (chaise):

If all you're looking for is an answer: (c+3)(3c-8)

OpenStudy (anonymous):

ac=-72, then rewrite c as 9c-8c 3c^2-8c+9c-24=c(3c-8)+3(3c-8)=(3c-8)(c+3)

OpenStudy (anonymous):

by rewriteing the middle terms using factors of 3*(-24)=-72 we can factor by grouping

OpenStudy (anonymous):

this process allows you to "reverse" the steps you would get from a FOIL if you started with (3c-8)(c+3)

OpenStudy (anonymous):

FOIL: (3c-8)(c+3) = 3c^2+9c -8c -24 F 0 I L